OpenStudy (anonymous):

Evaluate the line integral \[\int_c F\cdot dr\] where c is given by \[\hat r (t)=t\hat i +sint\hat j +cost \hat k\] \[0 \le t le \pi\] and: \[\hat r(t)=\hat i +cost \hat j -sint \hat k\] \[F(x, y, z)=z\hat i+y \hat j-x\hat k\] x=t y=sint z=cost \[\int_0^{\pi}(cost\hat i +sint\hat j-t\hat k)\cdot(\hat i+cos(t) \hat j -sint \hat k)dt\] \[\int_0^{\pi}(cost+sintcost+tsint)dt\] \[\int_0^{\pi} cost dt+\int_0^{\pi} sintcostdt+\int_0^{\pi}tsintdt\]

OpenStudy (anonymous):

\[sint]_0^{\pi}+\int_0^0 udu+[-tcost+sint]_0^{\pi}\]

OpenStudy (anonymous):

mukushla is not typing a reply…whats wrong with site...lol let me check it

OpenStudy (anonymous):

LOL...i think mukushla should type a reply

OpenStudy (anonymous):

its allright but what is that middle integral !!?

OpenStudy (anonymous):

u substitution u =sint du=cost dt sin(pi)=0 sin(0)=0

OpenStudy (anonymous):

almost done

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

cos Pi is -1 though

OpenStudy (anonymous):

so negative pi should be the answer?

OpenStudy (anonymous):

and cos (0) is 1 oh it cancels out!

OpenStudy (anonymous):

no that still doesn't seem right... :(

OpenStudy (anonymous):

i meant 0 for middle integral

OpenStudy (anonymous):

answer is \(\pi\) ?

OpenStudy (anonymous):

the first integral would be zero too and then we have left \[[-tcost+sint]_0^{\pi}=\pi\]

OpenStudy (anonymous):

YAY!

OpenStudy (anonymous):

We are DONE.....;-D